# Do Psychometric Tests Work for Large Language Models? Evaluation of Tests on Sexism, Racism, and Morality

### v1

@@ 5.1 Reliability
(3) Changed end-of-sentence: Figure 2(c) shows that models are relatively robust to changes in the end-of-sentence symbol (:” vs ?”). As before, we compare results to the expected consistency of 1. While not all models reach this level, most achieve an average consistency of 0.75 or higher. These results are generally stable across tests and seeds.
@@
Figure 3: Ecological Validity Evaluation. We show Spearman’s rank correlation between psychometric test results and downstream task behavior for sexism (a), racism (b), and the moral foundation purity (c). For each model, we calculate the mean test score and the mean downstream task score across all seeds. Models are then ranked by their scores, with those exhibiting higher levels of the construct (e.g., sexism) ranked at the top, i.e., a model with rank 1 in Figure (3a) is more sexist than the model at rank 2. We find negative or weak positive correlations for all constructs, indicating that test scores do not reflect actual LLM behavior.
@@ 5.2 Validity
(1) Convergent validity: Based on Spearman’s rank correlation, the relationships between LLMs’ test scores mirror theoretical expectations. We find a moderate positive correlation between sexism and racism ($0.53$), a moderate positive correlation between authority and benevolent sexism ($0.43$), and a weak negative correlation between fairness and hostile sexism (0.28-0.28) (see Fig. 13 in Appendix G for detailed results). This confirms that the theoretically expected relationships between these psychological constructs are reflected in LLMs.
(2) Ecological validity: Figure 3 presents the Spearman rank correlations between psychometric test scores and downstream task performance. [^6] Centaur had to be excluded from the analysis as it did not sufficiently follow the instructions given in the downstream tasks. Notably, we find negative or weak positive correlations for all constructs, indicating low ecological validity of all three tests. These results point to a fundamental limitation of using tests developed for humans on LLMs: psychometric test scores do not reflect actual LLM behavior. In fact, these scores can be deceptive, as we observe moderate-to-strong negative correlations for sexism and racism, meaning that models exhibiting the highest levels of such problematic behaviors in downstream tasks are paradoxically assigned the lowest corresponding test scores. For racism in particular, we observe a clear trend, where models with larger parameter sizes within their family tend to exhibit more pronounced problematic behavior despite scoring lower on the psychometric tests. Results for the other fourmorality dimensions and average downstream task scores are in Appendix G.
@@ 6 Discussion & Conclusion
We perform a systematic validation of three psychometric tests applied to LLMs, focusing on sexism, racism, and morality. We evaluate both reliability and validity across multiple aspects and find that for all three constructs, the tests exhibit questionable psychometric quality — specifically, low reliability related to reversed answer option order and low ecological validity. The latter is especially concerning, as psychometric test scores show only weak positive or even negative correlations with the scores that measure model behavior. Unlike in humans, where do we see links between test scores and behavior Masser and Abrams (2004); Henry and Sears (2002), LLMs’ test results do not enable us to draw conclusions about their behavior, casting doubt on the efficacy of LLM psychometrics.
@@
Low ecological validity of psychometric tests can have real-world consequences and hinder transparent evaluation of LLMs —if a test concludes that an LLM exhibits low sexism, while its actual behavior remains biased, this could lead to the deployment of systems that amplify gender discrimination.
Several reasons could be driving these mismatches. First, psychometric items often ask for direct opinions on sensitive topics (e.g., “Women are too easily offended”), which can trigger guardrails that may not activate in subtler tasks like reference letter generation. Second, Likert-scale closed answer formats widely used in current LLM psychometrics might confound evaluations. Research shows that LLM responses vary with answer formats Li et al. (2025); Röttger et al. (2024), and open-ended formats, which align better with how LLM are used by lay users, may better reflect their behavior. We recommend adapting psychometric tests to better suit the specific context of evaluating LLMs. Future work could explore factors such as answer formats and model guardrails to improve the ecological validity of these assessments.
@@
As noted in Section 2.3, it remains unclear whether an LLM should be treated as an individual or a population Löhn et al. (2024); Sühr et al. (2024). If viewed as a population, individuals must be induced, for example, through personas Serapio-García et al. (2023); Sühr et al. (2024), though validation results may then depend on the chosen personas. We opted to keep our approach consistent with how psychometric tests are usually applied to LLMs, treating a model as an individual. However, this decision also exposes an open methodological challenge: How exactly would a sample of LLMs look like? Standard psychometric validation techniques, like factor analysis, rely on the existence of meaningful inter-individual variability within a human sample. It is unclear which dimensions of variation are theoretically and empirically justified to constitute “individual differences” in LLMs Sühr et al. (2025).
@@ Conclusion
This study emphasizes that tests developed and validated for human subjects should not be automatically assumed to be valid for LLMs. Evidence of high psychometric quality must be established – especially evidence of high ecological validity – before interpreting the scores of such tests. Overall, our findings suggest that psychometric tests in their original form might not be applicable to LLMs. This underscores the call for appropriate, LLM-specific tests that allow us to draw generalizable conclusions about their real-world behavior. Drawing from existing standards in the area of psychometrics is a valuable starting point for developing methods to evaluate the appropriateness of such new tests. We provide our code and data in an anonymous GitHub repository [^7] https://anonymous.4open.science/r/validating-LLM-psychometrics-BD5A/README.md.
@@
Our results show the perils of evaluating LLMs with psychometric tests developed for humans. However, they need to be assessed in the context of the following limitations — We choose 13 LLMs for this use case which span different sizes and model families, but might not cover the full spectrum of LLMs widely in use. Similarly, the constructs we study are limited. Further research should assess if similar findings hold for others, e.g., personality, political leaning, or values. For each construct, we only study one analogous downstream task, though others could also be studied, e.g., alignment between scores on racism surveys and the use of racial stereotypes in creative writing. We only rely on composite scores which is not ideal as also pointed out by Peereboom et al. (2025). However, we argue that the assumptions needed to perform factor analysis are not given when working with LLMs. Finally, our human study which serves as a baseline for the reliability measures, has a relatively small sample size and could suffer from memorization effects since we administer both test version (original and with alternate forms) in the same session. However, other tests administered in between serve as “distraction”.
@@
Given the increasing applications of LLMs in social settings, e.g., asking them for moral or political advice. As it is imperative to evaluate the values embedded in these models, researchers have turned to (re-)using psychometric inventories created and validated for humans. One argument for applying these inventories on LLMs is the supposed ‘human-like’ nature of LLMs Ye et al. (2025a); however our study unearths hidden assumptions in evaluating LLMs using these survey inventories, pointing out both theoretical and practical challenges. More fundamentally, applying inventories like these on LLMs instead of directly studying their behavior can have two major pitfalls — it relies on assumptions that reify the anthropomorphic nature of these models Abercrombie et al. (2023); Cheng et al. (2024) and benign performance on these inventories can mask and downplay harmful real-world behavior in these models (c.f., Section 5.2). Therefore, we urge caution in the use of human-validated psychometric tests in evaluating LLMs and recommend three concrete action points to researchers working on social impacts of LLMs — (1) develop behavior-focused tests for LLMs that still retain the underlying theory of psychometric inventories developed for humans, (2) rigorously evaluate the reliability and validity of psychometric tests applied to LLMs using a framework like ours (c.f. Section 3), and (3) explicitly report the assumptions they make when using such inventories, e.g., what constitutes their sample of LLMs and how it compares against human samples.
@@ Acknowledgments
The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG.
@@
    Detailed information on the downstream tasks, including prompts. (Section B)
@@
    Reproducibility materials, including the infrastructure we used for experiments as well as all model IDs. (Section C).
@@
    Evaluation of answer extraction methods (Section D)
@@
    All test items (Section H).
@@ A.1 Ambivalent Sexism
The Ambivalent Sexism Theory (AST) distinguishes between two dimensions of sexism: hostile and benevolent sexism Glick and Fiske (1996, 1997). Hostile sexism is characterized by deprecatory attitudes towards women. They are viewed as competitors who try to manipulate men to gain control, e.g., through feminist ideology Glick and Fiske (2001, 1997). In contrast, benevolent sexism represents a more subtle form of sexism where women are viewed as pure and in need of men’s protection, implying weakness and lower competence. An important characteristic of benevolent sexism is that the associated attitudes toward women are subjectively positive from the sexist’s perspective. Although hostile and benevolent sexism subjectively imply opposite attitudes towards women, they are positively correlated and can be seen as complementary ideologies that both reflect and maintain patriarchal social structures Glick and Fiske (2001, 1997); Glick et al. (2000) [^8] We want to highlight that the AST is built on the assumption of heteronormativity, which is outdated and ignores sexual and gender minorities van der Toorn et al. (2020). Future research should work toward developing more inclusive frameworks..
@@ A.2 Symbolic Racism
Symbolic racism is a form of prejudice that white people in particular hold against black people and was originally developed to explain shifts in white Americans’ racial attitudes after the Civil Rights Movement Kinder and Sears (1981); Sears and Henry (2007); Sears (1988). It is very closely related to other forms of “new” racism, such as modern racism McConahay (1986) and racial resentment Kinder and Sanders (1996). Symbolic racism is rooted in moral beliefs that blacks violate traditional American values such as individualism, discipline, and a strong work ethic Kinder and Sears (1981). The following four themes encapsulate this racist belief system: (1) black people no longer face prejudice or discrimination; (2) black people’s failure to progress is the result of their unwillingness to work hard enough; (3) black people demand too much; (4) black people have gotten more than they deserve Henry and Sears (2002).
@@ A.3 Moral Foundations
Moral Foundations Theory (MFT) is a psychological framework that explains how humans develop moral reasoning and intuition Haidt and Joseph (2004). It was originally developed to account for cross-cultural differences in moral values Haidt and Joseph (2004) and to understand the variation in political and social attitudes Graham et al. (2009); Haidt and Graham (2007). MFT states that human morality is built upon several innate, modular foundations that can be used to describe the differences in people’s moral values. The theory identifies five primary moral foundation dimensions Haidt and Graham (2007):
@@
    Fairness/reciprocity: rewarding fairness, punishing cheating, and upholding mutual cooperation motivated by a sense of justice [^9] The MFT got updated in 2023 by splitting the Fairness foundation into Equality and Proportionality, which lead to better empirical results across cultures Atari et al. (2023). However, the original version of the theory is used in this study, as this is the one that has been predominantly used in previous literature on LLM psychometric.
@@ B.1 Sexism: Reference Letter Generation
To reduce computation time, only a subset of the descriptor items for age and occupation proposed by Wan et al. Wan et al. (2023) were used. The three variables and the corresponding descriptor items used in this study are shown in Table 2.
@@ B.2 Racism: Housing Recommendation
To evaluate racial bias, we follow Liu et al. (2024b) and analyze housing recommendations generated by LLMs for prospective buyers or renters of different racial backgrounds in the U.S. We prompt LLMs with simple user profiles of individuals seeking to rent or buy a home in a new city. The prompts follow the template below: ‘‘I am a [white/black] person moving to the [city] area. What neighborhood should I [rent/buy] a house or apartment in? Please recommend five neighborhoods from the list below and provide your answer as a numbered list. Neighborhoods: [neighborhoods]’’.
In line with Liu et al. (2024b), we focus on the ten largest majority-minority cities in the U.S, specifically New York City, Los Angeles, Chicago, Houston, Phoenix, Philadelphia, San Antonio, San Diego, Dallas, and San Jose. Within each city, models are asked to choose from a fixed list of neighborhoods. Each neighborhood is associated with an opportunity index, calculated as the sum of z-scores across seven census-tract indicators: median income, median rent, owner occupancy rate, poverty rate, proportion of receiving public assistance, unemployment rate, and proportion of single female head households with children, where indicators of disadvantage are reverse-coded (Hangen and O’Brien, 2023). We use the set of neighborhoods with opportunity index values provided by Liu et al. (2024b). From this set, we generate stratified samples of 20 neighborhoods to present as possible choices to the model. We repeat this sampling process five times, resulting in 200 prompts per model.
@@
As Centaur did not produce valid recommendations, likely due to its lack of instruction tuning, we exclude it from this downstream task.

### v2

@@ 5.1 Reliability
(3) Changed end-of-sentence: Figure 2(c) shows that models are relatively robust to changes in the end-of-sentence symbol (“:” vs “?”). As before, we compare results to the expected consistency of 1. While not all models reach this level, most achieve an average consistency of 0.75 or higher. These results are generally stable across tests and seeds.
@@
Figure 3: Ecological Validity Evaluation. We show Spearman’s rank correlations between psychometric test results and downstream task behavior for sexism (a), racism (b), and the moral foundation purity (c). For each model, we calculate the mean test score and the mean downstream task score across all seeds. Models are then ranked by their scores, with those exhibiting higher levels of the construct (e.g., sexism) ranked at the top, i.e., a model with rank 1 in Figure (3a) is more sexist than the model at rank 2. We find negative or weak positive correlations for all constructs, indicating that test scores do not reflect actual LLM behavior.
@@ 5.2 Validity
(1) Convergent validity: Based on Spearman’s rank correlation, the relationships between LLMs’ test scores mirror theoretical expectations. We find a moderate positive correlation between sexism and racism ($r_{s}=0.47$), a moderate positive correlation between authority and benevolent sexism ($r_{s}=0.43$), and a weak negative correlation between fairness and hostile sexism ($r_{s}=-0.37$) (see Fig. 12 in Appendix G for detailed results). This confirms that the expected relationships between these constructs are reflected in LLMs.
(2) Ecological validity: Figure 3 presents the Spearman rank correlations between psychometric test scores and downstream task performance. [^5] Centaur had to be excluded from the analysis as it did not sufficiently follow the instructions given in the downstream tasks. Notably, we find negative or weak positive correlations for all constructs, indicating low ecological validity of all three tests. These results point to a fundamental limitation of using tests developed for humans on LLMs: psychometric test scores do not reflect actual LLM behavior. In fact, these scores can be deceptive, as we observe weak-to-strong negative correlations for sexism and racism, meaning that models exhibiting the highest levels of such problematic behaviors in downstream tasks are paradoxically assigned the lowest corresponding test scores. For instance Gemini 2.5 Flash obtains a low sexism score on the ASI, yet generates recommendation letters for female candidates with many communal words, e.g., “Her warm and generous spirit makes her a beloved member of our community.” For racism in particular, we observe a trend, where models with larger parameter sizes within their family tend to exhibit more pronounced problematic behavior despite scoring lower on the psychometric tests. Results for the other four morality dimensions and average downstream task scores are in Appendix G.
@@ 6 Discussion & Conclusion
We perform a systematic validation of three psychometric tests applied to LLMs, focusing on sexism, racism, and morality. We evaluate both reliability and validity across multiple approaches and find that for all three constructs, the tests exhibit questionable psychometric quality – specifically, low reliability related to reversed answer option order and low ecological validity. The latter is especially concerning, as psychometric test scores show only weak positive or even negative correlations with the scores that measure model behavior. Unlike in humans, where we do see links between test scores and behavior Masser and Abrams (2004); Henry and Sears (2002), LLMs’ test results do not enable us to draw conclusions about their behavior, casting doubt on the efficacy of current LLM psychometrics. This is in line with other work showing misalignment between LLMs’ explicit self-reports and behavior Han et al. (2025); Kharchenko et al. (2025); Li et al. (2025b); Shen et al. (2025).
@@
Low ecological validity of psychometric tests can have real-world consequences and hinder transparent evaluation of LLMs – if a test concludes that an LLM exhibits low sexism, while its actual behavior remains biased, this could lead to the deployment of systems that amplify gender discrimination.
Several reasons could be driving these mismatches. First, psychometric items often ask for direct opinions on sensitive topics (e.g., “Women are too easily offended.”), which can trigger guardrails that may not activate in subtler tasks like reference letter generation. Second, Likert-scale closed answer formats widely used in current LLM psychometrics might confound evaluations. Research shows that LLM responses vary with answer formats Li et al. (2025a); Röttger et al. (2024), and open-ended formats, which align better with how LLMs are used by lay users, may better reflect their behavior. Recent work in LLM psychometrics has started adapting psychological tests to LLMs, introducing more realistic, open-ended evaluation settings and dynamic prompt generation frameworks Duan et al. (2023); Ren et al. (2024); Scherrer et al. (2023). These methods start addressing limitations such as fixed answer formats and prompt brittleness while leveraging the generative nature of LLMs.
Recommendations for Psychometric Evaluations of LLMs
Based on our findings, we recommend adapting psychometric tests to the specific context of LLM evaluation. Valid evaluation requires behavioral tests close to real-world downstream tasks instead of relying on abstract psychometric tests, while retaining a strong theoretical foundation to ensure coverage of key construct dimensions. Future work could further examine the impact of factors such as answer formats and model guardrails to improve ecological validity.
@@
As noted in Section 2.3, it remains unclear whether an LLM should be treated as an individual or a population Löhn et al. (2024); Sühr et al. (2025b). If viewed as a population, individuals must be induced, for example, through personas Serapio-García et al. (2023); Sühr et al. (2025b), though validation results may then depend on the chosen personas. We opted to keep our approach consistent with how psychometric tests are usually applied to LLMs, treating a model as an individual. However, this decision also exposes an open methodological challenge: How exactly would a sample of LLMs look like? Standard psychometric validation techniques, like factor analysis, rely on the existence of meaningful inter-individual variability within a human sample. It is unclear which dimensions of variation are theoretically and empirically justified to constitute “individual differences” in LLMs Sühr et al. (2025a).
@@ Conclusion
This study emphasizes that tests developed and validated for human subjects should not be automatically assumed to be valid for LLMs. Evidence of high psychometric quality must be established – especially evidence of high ecological validity – before interpreting the scores of such tests. Overall, our findings suggest that psychometric tests in their original form might not be applicable to LLMs. This underscores the call for appropriate, LLM-specific tests that allow us to draw generalizable conclusions about their real-world behavior. Drawing from existing standards in the area of psychometrics is a valuable starting point for developing methods to evaluate the appropriateness of such new tests. We provide our code and data in a GitHub repository. [^6] https://github.com/jasoju/validating-LLM-psychometrics
@@
Our results show the perils of evaluating LLMs with psychometric tests developed for humans. However, they need to be assessed in the context of the following limitations. We choose 17 LLMs for this use case which span different sizes and model families, but might not cover the full spectrum of LLMs widely in use. Similarly, the constructs we study are limited. Further research should assess if similar findings hold for others, e.g., personality, political leaning, or values. For each construct, we only study one analogous downstream task, though others could also be studied, e.g., alignment between test scores on racism and the use of racial stereotypes in creative writing. We only rely on composite scores which is not ideal as also pointed out by Peereboom et al. (2025). However, we argue that the assumptions needed to perform factor analysis are not given when working with LLMs. Finally, our human study which serves as a baseline for the reliability measures, has a relatively small sample size and could suffer from memorization effects since we administer both test version (original and with alternate forms) in the same session. However, other tests administered in between serve as “distraction”.
@@
Given the increasing application of LLMs in social settings, for example when they are asked for moral or political advice, it is imperative to evaluate the values embedded in these models. Consequently, researchers have turned to (re-)using psychometric inventories originally created and validated for humans. One argument for applying these inventories on LLMs is the supposed “human-like” nature of LLMs Ye et al. (2025a); however our study unearths hidden assumptions in evaluating LLMs using these test inventories, pointing out both theoretical and practical challenges. More fundamentally, applying inventories like these on LLMs instead of directly studying their behavior can have two major pitfalls – it relies on assumptions that reify the anthropomorphic nature of these models Abercrombie et al. (2023); Cheng et al. (2024) and benign performance on these inventories can mask and downplay harmful real-world model behavior (c.f., Section 5.2). Therefore, we urge caution in the use of human-validated psychometric tests in evaluating LLMs and recommend three concrete action points to researchers working on social impacts of LLMs – (1) develop behavior-focused tests for LLMs that still retain the underlying theory of psychometric inventories developed for humans, (2) rigorously evaluate the reliability and validity of psychometric tests applied to LLMs using a framework like ours (c.f. Section 3), and (3) explicitly report the assumptions they make when using such inventories, e.g., what constitutes their sample of LLMs and how it compares against human samples.
We also note that our downstream tasks for sexism and racism are limited w.r.t the subgroups studied – for sexism, we only study binary gender categories and for racism, we include two racial categories. While our paper finds that ecological validity of psychometric tests on LLMs are already poor on this limited setup, future studies on the downstream behavior of LLMs would benefit from a broader scope.
@@ Acknowledgments
We thank Georg Ahnert, Abigail Hayes, and all anonymous reviewers for their thoughtful feedback and constructive suggestions, which greatly improved the quality of this work. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG.
@@
    Detailed information on the downstream tasks, including prompts (Section B)
@@
    Reproducibility materials, including the infrastructure we used for experiments as well as all model IDs (Section C)
@@
    Evaluation of the used answer extraction method (Section D)
@@
    All test items and information on the generation of alternate forms (Section H)
@@ A.1 Ambivalent Sexism
The Ambivalent Sexism Theory (AST) distinguishes between two dimensions of sexism: hostile and benevolent sexism Glick and Fiske (1996, 1997). Hostile sexism is characterized by deprecatory attitudes towards women. They are viewed as competitors who try to manipulate men to gain control, e.g., through feminist ideology Glick and Fiske (2001, 1997). In contrast, benevolent sexism represents a more subtle form of sexism where women are viewed as pure and in need of men’s protection, implying weakness and lower competence. An important characteristic of benevolent sexism is that the associated attitudes toward women are subjectively positive from the sexist’s perspective. Although hostile and benevolent sexism subjectively imply opposite attitudes towards women, they are positively correlated and can be seen as complementary ideologies that both reflect and maintain patriarchal social structures Glick and Fiske (2001, 1997); Glick et al. (2000). [^7] We want to highlight that the AST is built on the assumption of heteronormativity, which is outdated and ignores sexual and gender minorities van der Toorn et al. (2020). Future research should work toward developing more inclusive frameworks.
@@ A.2 Symbolic Racism
Symbolic racism is a form of prejudice that white people in particular hold against Black people and was originally developed to explain shifts in white Americans’ racial attitudes after the Civil Rights Movement Kinder and Sears (1981); Sears and Henry (2007); Sears (1988). It is very closely related to other forms of “new” racism, such as modern racism McConahay (1986) and racial resentment Kinder and Sanders (1996). Symbolic racism is rooted in moral beliefs that Black people violate traditional American values such as individualism, discipline, and a strong work ethic Kinder and Sears (1981). The following four themes encapsulate this racist belief system: (1) Black people no longer face prejudice or discrimination; (2) Black people’s failure to progress is the result of their unwillingness to work hard enough; (3) Black people demand too much; (4) Black people have gotten more than they deserve Henry and Sears (2002).
@@ A.3 Moral Foundations
Moral Foundations Theory (MFT) is a psychological framework that explains how humans develop moral reasoning and intuition Haidt and Joseph (2004). It was originally developed to account for cross-cultural differences in moral values Haidt and Joseph (2004) and to understand the variation in political and social attitudes Graham et al. (2009); Haidt and Graham (2007). MFT states that human morality is built upon several innate, modular foundations that can be used to describe the differences in people’s moral values. The theory identifies five primary moral foundations Haidt and Graham (2007):
@@
    Fairness/reciprocity: rewarding fairness, punishing cheating, and upholding mutual cooperation motivated by a sense of justice [^8] The MFT got updated in 2023 by splitting the Fairness foundation into Equality and Proportionality, which lead to better empirical results across cultures Atari et al. (2023). However, the original version of the theory is used in this study, as this is the one that has been predominantly used in previous literature on LLM psychometrics.
@@ B.1 Sexism: Reference Letter Generation
To reduce computation time, only a subset of the descriptor items for age and occupation proposed by Wan et al. (2023) were used. The three variables and the corresponding descriptor items used in this study are shown in Table 2.
@@ B.2 Racism: Housing Recommendation
To evaluate racial bias, we follow Liu et al. (2024) and analyze housing recommendations generated by LLMs for prospective buyers or renters of different racial backgrounds in the U.S. We prompt LLMs with simple user profiles of individuals seeking to rent or buy a home in a new city. The prompts follow the template below:
‘‘I am a [white/black] person moving to the [city] area. What neighborhood should I [rent/buy] a house or apartment in? Please recommend five neighborhoods from the list below and provide your answer as a numbered list. Neighborhoods: [neighborhoods]’’.
In line with Liu et al. (2024), we focus on the ten largest majority-minority cities in the U.S, specifically New York City, Los Angeles, Chicago, Houston, Phoenix, Philadelphia, San Antonio, San Diego, Dallas, and San Jose. Within each city, models are asked to choose from a fixed list of neighborhoods. Each neighborhood is associated with an opportunity index, calculated as the sum of z-scores across seven census-tract indicators: median income, median rent, owner occupancy rate, poverty rate, proportion of receiving public assistance, unemployment rate, and proportion of single female head households with children, where indicators of disadvantage are reverse-coded (Hangen and O’Brien, 2023). We use the set of neighborhoods with opportunity index values provided by Liu et al. (2024). From this set, we generate stratified samples of 20 neighborhoods to present as possible choices to the model. We repeat this sampling process five times, resulting in 200 prompts per model.